OpenStudy (anonymous):

I have 3 more questions then i dont need help anymore for algebra 1. Fan and medal!! just wait for me to post the equation

OpenStudy (anonymous):

what is the value of sin(84degrees) to the nearest one thousandth

OpenStudy (anonymous):

1

OpenStudy (anonymous):

without a calculator?

OpenStudy (anonymous):

what is the value of cos 34 degrees to the nearest ten thousandth

OpenStudy (anonymous):

-.84

OpenStudy (anonymous):

.829

OpenStudy (anonymous):

what is the answer to the sin questoin because 1 isnt the right answer but hwtrain got the second one right

OpenStudy (anonymous):

.995

OpenStudy (anonymous):

one more question

OpenStudy (anonymous):

@Darealest123 -.84 is correct for cos(34) and sin(84) is .73

OpenStudy (anonymous):

@Darealest123 pass the medal to me

OpenStudy (anonymous):

simplify the rational expression by rationalizing the denominator \[\frac{ 4 \sqrt{150} }{ \sqrt{189x} }\]

OpenStudy (anonymous):

How is that possible think of unit circle

OpenStudy (anonymous):

sorry kat the answers you gave werent even options for the question

OpenStudy (anonymous):